OpenStudy (anonymous):

On an air track, a 400.0g glider moving to the right at 1.70m/s collides elastically with a 497.0g glider moving in the opposite direction at 3.25m/s Find the velocity of first glider after the collision. m/s to the left Find the velocity of second glider after the collision. m/s to the right

OpenStudy (anonymous):

 you must know that if the collision is elastic then three things remains constant before and after collision:- 1) sum of Kinetic Energy's 2) sum of momentums 3) total energy let us use 2nd one here momentum=MV let us take initial condition

OpenStudy (anonymous):



OpenStudy (anonymous):

MU-mu (as opp. direction)=MV+mv 0.93525=MV+mv------------------(a) by conservation of Kinetic energy MU^2+mu^2=MV^2+mv^2 4.0935625=(MV)^2/M+(mv)^2/m----------------(b) two equations two unknowns just solve them and get answer:)